### v2

# Fullness of $q$-Araki-Woods factors

Abstract.
The $q$-Araki-Woods factor associated to a group of orthogonal transformations on a real separable Hilbert space $\mathsf{H}_{\mathbb{R}}$ is full as soon as $\dim\mathsf{H}_{\mathbb{R}}\geq 2$.
Key words and phrases:
$q$-Araki-Woods factors; full factors; type III factors
2020 Mathematics Subject Classification:
Primary: 46L36; Secondary 46L10, 46L53, 46L65
1. Introduction
A von Neumann factor $\mathsf{M}$ is called full if, for any bounded net ${x_{i}}$ in $\mathsf{M}$ such that $\lim_{i}\|\varphi(\cdot x_{i})-\varphi(x_{i}\cdot)\|_{\mathsf{M}_{*}}=0$ for all $\varphi\in\mathsf{M}_{*}$, there exists a bounded net ${\lambda_{i}}$ of scalars such that $\lim_{i}(x_{i}-\lambda_{i})=0$-strongly [Con1]. Here, $\mathsf{M}_{*}$ denotes the predual of $\mathsf{M}$ i.e. the space of normal linear functionals on $\mathsf{M}$. A full factor is always non-injective or non-hyperfinite i.e. it cannot be approximated by finite-dimensional subalgebras. For separable type II1 factors, the negation of fullness is equivalent to Murray and von Neumann’s notion of Property Gamma [Con2]. Connes established several other equivalent conditions for fullness in type II1 factors in his seminal work [Con2] on the classification program of injective factors. Some of these equivalent conditions, such as the spectral gap property, do not seem to generalize easily to type III factors. However, in a major breakthrough recently, Marrakchi [Mar] proved that fullness for separable type III1 factors is equivalent to the assertion that the $C^{*}$-algebra generated by $\mathsf{M}$ and its commutant $\mathsf{M}^{\prime}$, represented in the standard form, contains compact operators. The focus of this paper is to utilize one of these equivalent conditions to establish, in full generality, that the $q$-Araki-Woods factors are full.
Recall that the $q$-Araki-Woods factors, denoted as $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$, are associated with a number $q\in(-1,1)$ and a strongly continuous one-parameter group of orthogonal transformations $\{U_{t}\}_{t\in\mathbb{R}}$ on a real Hilbert space $\mathsf{H}_{\mathbb{R}}$ (see the construction in Section 2). These von Neumann algebras were introduced by Hiai [Hia], combining the quasi-free deformations of Shlyakhtenko [Shl] and $q$-deformations of Bożejko-Speicher [BS]. While various structural properties are known for the free Araki-Woods factors $\Gamma_{0}(\mathsf{H}_{\mathbb{R}},U_{t})$, such as fullness, solidity, strong solidity, the absence of Cartan subalgebras, and complete metric approximation properties [Shl, HRau, HRic, BHV], many of these properties remain elusive for non-zero $q$, with only a few of the relevant results available [ABW, BM, BMRW, KSW, Nou, SW]. For instance, the fundamental question of factoriality of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ was only recently addressed in complete generality in [KSW], with some major success achieved earlier in [BMRW]. Furthermore, the factors $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ are known to be non-injective [KSW, Nou]. Drawing inspiration from the free/undeformed case, it is then natural to ask whether $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is full as well for non-zero $q$. This problem has been considered earlier in [Sni, HI, KSW] with results obtained in various specific cases. In this paper, we address the remaining unresolved cases and hence present the following result in full generality.
Theorem 1.1.
Let $q\in(-1,1)$, let $\mathsf{H}_{\mathbb{R}}$ be a real separable Hilbert space with $\dim\mathsf{H}_{\mathbb{R}}\geq 2$, and let $\{U_{t}\}_{t\in\mathbb{R}}$ be a strongly continuous group of orthogonal transformations on $\mathsf{H}_{\mathbb{R}}$. Then the $q$-Araki-Woods factor $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is full.
We provide a concise overview of previous attempts and highlight our contribution to the proof. When the representation $\{U_{t}\}_{t\in\mathbb{R}}$ is trivial and $q\neq 0$, Śniady [Sni] proved the fullness of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}})$ for a large-dimensional real Hilbert space $\mathsf{H}_{\mathbb{R}}$, dependent on the value of $q$ (including the case when $\dim\mathsf{H}_{\mathbb{R}}=\infty$); additionally, Miyagawa-Speicher [MS] established the fullness of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}})$ for a finite-dimensional Hilbert space $\mathsf{H}_{\mathbb{R}}$ with $\dim\mathsf{H}_{\mathbb{R}}\geq 2$. On the other hand, there are mainly two papers, [KSW] and [HI], which have dealt with the fullness of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ for $q\neq 0$ and $\{U_{t}\}_{t\in\mathbb{R}}$ non-trivial. In order to describe them, it is crucial to first mention that the Hilbert space $\mathsf{H}_{\mathbb{R}}$ can be decomposed into direct sums of almost periodic $\mathsf{H}_{\mathbb{R}}^{ap}$ and weakly mixing $\mathsf{H}_{\mathbb{R}}^{wm}$ subspaces, where $\mathsf{H}_{\mathbb{R}}^{ap}$ is the direct sum of all finite-dimensional invariant subspaces of $\{U_{t}\}_{t\in\mathbb{R}}$, and $\mathsf{H}_{\mathbb{R}}^{wm}$ encompasses the rest. When $\mathsf{H}_{\mathbb{R}}$ is finite-dimensional with $\dim\mathsf{H}_{\mathbb{R}}\geq 2$, $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is shown to be full in [KSW]. The proof involves establishing the non-injectivity of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ by combining results on conjugate variables and general findings of Nelson [Nel]. It then proceeds to demonstrate Ozawa’s AO property for $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ before invoking the general results of Houdayer-Raum [HRau] to conclude fullness. In contrast, Houdayer and Isono [HI] show that if $\mathsf{H}_{\mathbb{R}}^{wm}\neq 0$ or if the point spectrum, the set of of eigenvalues of $A$, is infinite and bounded, then $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is full. Here, $A$ denotes the analytic generator of the one-parameter group of unitaries $\{U_{t}\}_{t\in\mathbb{R}}$, extended linearly to the complexification of $\mathsf{H}_{\mathbb{R}}$. Their proof follows the ideas of [Sni] from the tracial case, i.e. when $\{U_{t}\}_{t\in\mathbb{R}}$ is trivial. However, their approach encounters difficulties when there is a sequence of eigenvalues growing sufficiently fast to infinity. In this paper, we overcome such challenges by adopting a different approach.
Our arguments involve a meticulous evaluation of certain Wick products in the centralizer of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ associated with vectors in a $2\times 2$ block of $\mathsf{H}_{\mathbb{R}}$. As in [BMRW], we make a crucial use of the non-triviality of the group $\{U_{t}\}_{t\in\mathbb{R}}$; however, it is worth noting that the methods in [BMRW] only work for the weak convergence of the concerning sequences, which does not apply to the study of fullness. In our approach, we establish several norm convergence results involving the Wick products and creation/annihilation operators. The key novelty in these convergence results lies in the utilization of an ‘averaging technique’, where one will observe that the sequences themselves may fail to converge. For instance, when $\xi$ and $\eta$ are two unit vectors in $\mathsf{H}_{\mathbb{R}}$, the bounded sequence $T_{n}=(1-q)^{n}c(\xi)^{n}c(\eta)^{n}$ never converges in norm, although it does converge to 00 in weak operator topology, where $c(\xi)$ denotes the creation operator on the (deformed) Fock space. Nonetheless, we demonstrate that the Cesáro sum of $T_{n}$ converges to 00 in norm with the additional assumption that $\xi,\eta$ are orthogonal to each other (see Proposition 3.4). The main idea then is to carefully select such vectors $\xi$ and $\eta$, and employ these convergence results to prove fullness (see Theorem 1.1). It is crucial to emphasize that the methods employed in this paper are not applicable to the cases considered in the previous two attempts [KSW, HI]. Thus, a combination of both [KSW] and [HI] is necessary to conclusively establish Theorem 1.1 in full generality.
This paper is organized as follows. We give a brief overview of the construction of $q$-Araki-Woods factors in Section 2. Section 3 is the core of the article where we carry our main computations on various convergence results. Finally, we put all our efforts together in Section 4 in proving the main result as mentioned above.
2. Preliminaries on $q$-Araki-Woods factors
We briefly describe construction of the $q$-Araki-Woods factors and refer the readers to [Hia] for a detailed treatment. As a convention, all Hilbert spaces are separable and inner products are linear in the second variable. For any real or complex Hilbert space $H$, ${B}(H)$ denotes the algebra of bounded linear operators on $H$.
Let $\mathsf{H}_{\mathbb{R}}$ be a real Hilbert space, and let $\{U_{t}\}_{t\in\mathbb{R}}$ be a strongly continuous orthogonal representation of $\mathbb{R}$ on $\mathsf{H}_{\mathbb{R}}$. Consider the complexification $\mathsf{H}_{\mathbb{C}}=\mathsf{H}_{\mathbb{R}}\otimes_{\mathbb{R}}\mathbb{C}$ of $\mathsf{H}_{\mathbb{R}}$, whose inner product is denoted by $\langle\cdot,\cdot\rangle_{\mathsf{H}_{\mathbb{C}}}$. Identify $\mathsf{H}_{\mathbb{R}}$ inside $\mathsf{H}_{\mathbb{C}}$ by $\mathsf{H}_{\mathbb{R}}\otimes 1$, so that $\mathsf{H}_{\mathbb{C}}=\mathsf{H}_{\mathbb{R}}+i\mathsf{H}_{\mathbb{R}}$. Given $\xi=\xi_{1}+i\xi_{2}\in\mathsf{H}_{\mathbb{C}}$ for some $\xi_{1},\xi_{2}\in\mathsf{H}_{\mathbb{R}}$, write the canonical involution $\bar{\xi}$ as
$\bar{\xi}=\xi_{1}-i\xi_{2}.$
We extend $\{U_{t}\}_{t\in\mathbb{R}}$ to a strongly continuous one parameter group of unitaries on the complexification $\mathsf{H}_{\mathbb{C}}$, which is still denoted by $\{U_{t}\}_{t\in\mathbb{R}}$. Let $A$ denote the positive, self-adjoint, and non-singular (unbounded) operator on $\mathsf{H}_{\mathbb{C}}$ obtained via Stone’s theorem, which satisfies $U_{t}=A^{it}$ for all $t\in\mathbb{R}$. Note that $\overline{A\xi}=A^{-1}\bar{\xi}$ for any $\xi$ in the domain of $A$. In particular, $\sigma(A)\cap(0,\infty)$ is symmetric around $1$ i.e. for $\lambda\neq 0$, $\lambda\in\sigma(A)$ if and only if $1/\lambda\in\sigma(A)$, where $\sigma(A)$ is the spectrum of $A$.
Following Shlyakhtenko [Shl], consider the deformed inner product $\langle\cdot,\cdot\rangle_{U}$ on $\mathsf{H}_{\mathbb{C}}$ defined by
$\langle\xi,\zeta\rangle_{U}=\langle 2(1+A^{-1})^{-1}\xi,\zeta\rangle\quad\forall\xi,\zeta\in\mathsf{H}_{\mathbb{C}}.$
We denote the completion of $\mathsf{H}_{\mathbb{C}}$ with respect to $\langle\cdot,\cdot\rangle_{U}$ by $\mathsf{H}_{U}$. The space $(\mathsf{H}_{\mathbb{R}},\|\cdot\|_{\mathsf{H}_{\mathbb{R}}})$ is naturally identified as a subspace of $\mathsf{H}_{U}$ such that $\mathsf{H}_{\mathbb{R}}+i\mathsf{H}_{\mathbb{R}}$ is dense in $\mathsf{H}_{U}$ and $\mathsf{H}_{\mathbb{R}}\cap i\mathsf{H}_{\mathbb{R}}=\{0\}$. For $q\in(-1,1)$, consider the $q$-Fock space $\mathcal{F}_{q}(\mathsf{H}_{U})$ given by
$\mathcal{F}_{q}(\mathsf{H}_{U})=\mathbb{C}\Omega\oplus\bigoplus_{n\geq 1}\mathsf{H}_{U}^{\otimes_{q}^{n}},$
where $\Omega$ is a distinguished unit vector in $\mathbb{C}$ (called vacuum vector) and $\mathsf{H}_{U}^{\otimes_{q}^{n}}$ is the completion of the algebraic tensor product $\mathsf{H}^{\otimes_{\text{alg}}^{n}}$ equipped with the following inner product:
$\langle\xi_{1}\otimes\ldots\otimes\xi_{n},\zeta_{1}\otimes\ldots\otimes\zeta_{n}\rangle_{q}=\sum_{\pi\in S_{n}}q^{|\pi|}\prod_{i=1}^{n}\langle\xi_{i},\zeta_{\pi(i)}\rangle_{U}$
for all $\xi_{i},\zeta_{i}\in\mathsf{H}_{U}$, $1\leq i\leq n$. Here $S_{n}$ denotes the symmetric group on $n$ elements and $|\pi|$ denotes the number of inversions for the permutation $\pi\in S_{n}$. That the above assignment yields a genuine inner product was proved in [BS]. On $\mathcal{F}_{q}(\mathsf{H}_{U})$, we define for each $\xi\in\mathsf{H}_{U}$ operators $c(\xi)$ (called left q-creation) and $c^{*}(\xi)$ (called left q-annihilation) by the following assignments:
$c(\xi)\Omega=\xi,$
$c(\xi)(\xi_{1}\otimes\ldots\otimes\xi_{n})=\xi\otimes\xi_{1}\otimes\ldots\otimes\xi_{n},\;\;\;\mbox{ and }$
$c^{*}(\xi)\Omega=0,$
$c^{*}(\xi)(\xi_{1}\otimes\ldots\otimes\xi_{n})=\sum_{i=1}^{n}q^{i-1}\langle\xi,\xi_{i}\rangle\xi_{1}\otimes\ldots\otimes\hat{\xi_{i}}\otimes\ldots\otimes\xi_{n},$
for all $\xi_{1}\otimes\ldots\otimes\xi_{n}\in\mathsf{H}_{U}^{\otimes_{q}^{n}}$. Here $\hat{\xi_{i}}$ notates the omitted $i$-th letter from the tensor. That $c(\xi)$ and $c^{*}(\xi)$ extend to bounded linear operators on $\mathcal{F}_{q}(\mathsf{H}_{U})$ being adjoint to each other is a well known fact (see [BS]). In a similar fashion, one defines the right $q$-creation and right $q$-annihilation operators $r(\xi)$ and $r^{*}(\xi)$ respectively for all $\xi\in\mathsf{H}_{U}$ as follows:
$r(\xi)\Omega=\xi,$
$r(\xi)(\xi_{1}\otimes\ldots\otimes\xi_{n})=\xi_{1}\otimes\ldots\otimes\xi_{n}\otimes\xi,\;\;\;\mbox{ and }$
$r^{*}(\xi)\Omega=0,$
$r^{*}(\xi)(\xi_{1}\otimes\ldots\otimes\xi_{n})=\sum_{i=1}^{n}q^{n-i}\langle\xi,\xi_{i}\rangle\xi_{1}\otimes\ldots\otimes\hat{\xi_{i}}\otimes\ldots\otimes\xi_{n},$
for all $\xi_{1}\otimes\ldots\otimes\xi_{n}\in\mathsf{H}_{U}^{\otimes_{q}^{n}}$. Again $r(\xi)$ and $r^{*}(\xi)$ are bounded operators which are adjoint to each other. The creation and annihilation operators satisfy the following $q$-commutation relations:
(1) $c^{*}(\xi)c(\eta)-qc(\eta)c^{*}(\xi)=\langle\xi,\eta\rangle_{U}$
$r^{*}(\xi)r(\eta)-qr(\eta)r^{*}(\xi)=\langle\xi,\eta\rangle_{U}$
for any $\xi,\eta\in\mathsf{H}_{U}$. We shall repeatedly use the above commutation relations without referring to them. In particular, the most useful use of these relations will be the following consequence:
$c^{*}(\xi)^{m}c(\eta)^{n}=q^{mn}c(\eta)^{n}c^{*}(\xi)^{m}$
whenever $\xi,\eta\in\mathsf{H}_{U}$ with $\langle\xi,\eta\rangle_{U}=0$ and $m,n\geq 1$.
We now consider the von Neumann algebra introduced by Hiai [Hia] (following Shlyakhtenko [Shl] and Bożejko-Speicher [BS]), denoted $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$, to be the von Neumann subalgebra of ${B}(\mathcal{F}_{q}(\mathsf{H}_{U}))$ generated by the set $\{W(\xi);\xi\in\mathsf{H}_{\mathbb{R}}\}$, where
$W(\xi)=c({\xi})+c^{*}(\xi),\;\;\;\;\;\;\;\xi\in\mathsf{H}_{\mathbb{R}}.$
The algebra $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is called the $q$-deformed Araki-Woods von Neumann algebra. If the representation $\{U_{t}\}_{t\in\mathbb{R}}$ is clear from the context, we shall denote the algebra $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ simply by $\mathsf{M}_{q}$. We shall always assume that $\dim\mathsf{H}_{\mathbb{R}}\geq 2$. Since it is now known that $\mathsf{M}_{q}$ is a factor i.e. $\mathsf{M}_{q}$ has trivial center (see [KSW, Hia]), we will call it a $q$-Araki-Woods factor.
The vacuum vector $\Omega$ is cyclic and separating for the factor $\mathsf{M}_{q}$, and hence induces a faithful normal state: $\varphi(\cdot)=\langle\Omega,\cdot\;\Omega\rangle_{q}$ on $\mathsf{M}_{q}$, called the $q$-quasi free state. It is immediate to verify that the GNS Hilbert space associated to $\varphi$ is $\mathcal{F}_{q}(\mathsf{H}_{U})$. Since $\Omega$ is a separating vector for both $\mathsf{M}_{q}$ and $\mathsf{M}_{q}^{\prime}$, there exist for each $\xi\in\mathsf{M}_{q}\Omega$ and $\zeta\in\mathsf{M}_{q}^{\prime}\Omega$ unique operators $x_{\xi}\in\mathsf{M}_{q}$ and $x^{\prime}_{\zeta}\in\mathsf{M}_{q}^{\prime}$ such that $\xi=x_{\xi}\Omega$ and $\zeta=x_{\zeta}^{\prime}\Omega$. Write
$W(\xi)=x_{\xi},\;\mbox{ and }\;W_{r}(\zeta)=x^{\prime}_{\zeta}.$
Note that the notation $W(\xi)$ is in line with the one introduced above for $\xi\in\mathsf{H}_{\mathbb{R}}$. The operators $W(\xi)$ and $W_{r}(\zeta)$ are respectively called left and right Wick operators. It is known that $\xi_{1}\otimes\ldots\otimes\xi_{n}\in\mathsf{M}_{q}\Omega$ for all $\xi_{i}\in\mathsf{H}_{\mathbb{C}},1\leq i\leq n$ and $\zeta_{1}\otimes\ldots\otimes\zeta_{m}\in\mathsf{M}_{q}^{\prime}\Omega$ for all $\zeta_{j}\in\mathsf{H}_{\mathbb{C}}\cap\operatorname{Dom}(A^{-1/2}),1\leq j\leq m$. By definition, the operators ${W}(\xi)$ and $W_{r}(\zeta)$ commute for all $\xi\in\mathsf{M}_{q}\Omega$ and $\zeta\in\mathsf{M}_{q}^{\prime}\Omega$. In fact, the commutant of $\mathsf{M}_{q}$ in $B(\mathcal{F}_{q}(\mathsf{H}_{U}))$ is given by
$\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})^{\prime}=\{W_{r}(\eta);\eta\in\mathsf{H}_{\mathbb{R}}^{\prime}\}^{{}^{\prime\prime}}$
where $\mathsf{H}_{\mathbb{R}}^{\prime}=\{\eta\in\mathsf{H}_{\mathbb{C}};\langle\xi,\eta\rangle_{U}\in\mathbb{R}\mbox{ for all }\xi\in\mathsf{H}_{\mathbb{R}}\}$. We have $\mathsf{H}_{\mathbb{C}}=\mathsf{H}_{\mathbb{R}}^{\prime}+i\mathsf{H}_{\mathbb{R}}^{\prime}$ and $\mathsf{H}_{\mathbb{R}}\cap i\mathsf{H}_{\mathbb{R}}=\{0\}$. It is easy to verify that
$W(\xi)=c({\xi})+c^{*}({\bar{\xi}})\;\;\mbox{ and }\;\;W_{r}(\zeta)=r(\zeta)+r^{*}(\bar{\zeta}^{r})$
for any $\xi\in\mathsf{H}_{\mathbb{C}}$ and $\zeta\in\mathsf{H}_{\mathbb{C}}\cap\operatorname{Dom}(A^{-1/2})$, where $\bar{\zeta}^{r}$ denotes the real part of $\zeta$ with respect to the decomposition $\mathsf{H}_{\mathbb{C}}=\mathsf{H}_{\mathbb{R}}^{\prime}+i\mathsf{H}_{\mathbb{R}}^{\prime}$. The formula for the left and right Wick operators are given as follows:
Proposition 2.1.
For any vectors $\xi_{1},\ldots,\xi_{n}\in\mathsf{H}_{\mathbb{R}}+i\mathsf{H}_{\mathbb{R}}$, the Wick-formula for operators in $\mathsf{M}_{q}$ is given by
$W(\xi_{1}\otimes\cdots\otimes\xi_{n})$ $=\sum_{i=0}^{n}\sum_{\sigma\in S_{n,i}}q^{|\sigma|}c(\xi_{\sigma(1)})\cdots c(\xi_{\sigma(i)})c^{*}(\bar{\xi}_{\sigma(i+1)})\cdots c^{*}(\bar{\xi}_{\sigma(n)})$
where $S_{n,j}$ is the set of those permutations in $S_{n}$ which are increasing on ${\{1,\ldots,j\}}$ and on $\{j+1,\ldots,n\}$.
Proposition 2.2.
For any vectors $\xi_{1},\ldots,\xi_{n}\in\mathsf{H}_{\mathbb{R}}^{\prime}+i\mathsf{H}_{\mathbb{R}}^{\prime}$, the Wick-formula for operators in $\mathsf{M}_{q}^{\prime}$ is given by
$W_{r}(\xi_{n}\otimes\cdots\otimes\xi_{1})$ $=\sum_{i=0}^{n}\sum_{\sigma\in S_{n,i}}q^{|\operatorname{flip}\circ\sigma|}r(\xi_{\sigma(1)})\cdots r(\xi_{\sigma(n-i)})r^{*}(\bar{\xi}_{\sigma(n-i+1)}^{r})\cdots r^{*}(\bar{\xi}_{\sigma(n)}^{r})$
$=\sum_{i=0}^{n}\sum_{\sigma\in S_{n,n-i}}q^{|\sigma|}r(\xi_{n+1-\sigma(n)})\cdots r(\xi_{n+1-\sigma(i+1)})r^{*}(\bar{\xi}_{n+1-\sigma(i)}^{r})\cdots r^{*}(\bar{\xi}_{n+1-\sigma(1)}^{r})$
where $S_{n,j}$ is the set of those permutations in $S_{n}$ which are increasing on ${\{1,\ldots,j\}}$ and on $\{j+1,\ldots,n\}$.
We adopt the following notations for our convenience.
  1. (1)
    $\xi_{1}\ldots\xi_{n}=\xi_{1}\otimes\ldots\otimes\xi_{n}$, for $\xi_{i}\in\mathsf{H}_{U}$, $1\leq i\leq n$.
  2. (2)
    $\xi^{n}=\xi^{\otimes n}$, $n\geq 1$ and $\xi^{0}=\Omega$ for all $\xi\in\mathsf{H}_{U}$.
  3. (3)
    $d_{0}=1$, $d_{n}=\prod_{j=1}^{n}(1-q^{j})$ for $n\geq 1$ and $d_{\infty}=\prod_{j=1}^{\infty}(1-q^{j})$
  4. (4)
    $C_{q}=\prod_{j=1}^{\infty}\frac{1}{1-|q|^{j}}$
  5. (5)
    $[n]_{q}=\sum_{j=0}^{n-1}q^{j}=\frac{1-q^{n}}{1-q}$
  6. (6)
    $[n]_{q}!=\prod_{j=1}^{n}[j]_{q}=\frac{d_{n}}{(1-q)^{n}}$
  7. (7)
    $\begin{pmatrix}n\\ k\end{pmatrix}_{q}=\frac{[n]_{q}!}{[k]_{q}![n-k]_{q}!}$ for any $1\leq k\leq n$.
The following are well known inequalities (see [BS]), which we shall use very frequently throughout the paper:
  1. (1)
    $d_{n}\leq C_{q}$,
  2. (2)
    $\frac{1}{d_{n}}\leq C_{q}$.
  3. (3)
    $\|c(\xi)^{n}\|\leq\sqrt{C_{q}}\|\xi^{n}\|_{q}$ for all $\xi\in\mathsf{H}_{U}$.
The following statements are either well-known or easy to verify.
Lemma 2.3.
Given $f\in\mathsf{H}_{U}$, and $\xi\in\mathsf{H}_{U}^{\otimes_{q}^{m}},\eta\in\mathsf{H}_{U}^{\otimes_{q}^{n}}$ for some $m,n\geq 1$, we have
$c^{*}(f)(\xi\otimes\eta)=(c^{*}(f)\xi)\otimes\eta+q^{m}\xi\otimes c^{*}(f)\eta$
Lemma 2.4.
For any $\xi\in\mathsf{H}_{U}$ and $n\geq 1$, we have $\|c(\xi)^{n}\|\leq C_{q}(1-q)^{-\frac{n}{2}}\|\xi\|^{\frac{n}{2}}.$
Proof.
This follows by recalling that $\|c(\xi)^{n}\|\leq\sqrt{C_{q}}\|\xi^{n}\|_{q}$, as well as by noting that $[n]_{q}!=d_{n}(1-q)^{-n}\leq C_{q}(1-q)^{-n}$ for all $n\geq 1$. ∎
Corollary 2.5.
For any $\xi\in\mathsf{H}_{U}$, the sequences
  1. (1)
    $(1-q)^{\frac{n}{2}}\|\xi\|_{U}^{-n}c(\xi)^{n}$
  2. (2)
    $(1-q)^{\frac{n}{2}}\|\xi\|_{U}^{-n}r(\xi)^{n}$
are bounded, both with an upper bound given by $C_{q}$.
We end this section by recalling the general structure of a one parameter group of orthogonal operators. Let $\{U_{t}\}_{t\in\mathbb{R}}$ be a strongly continuous orthogonal representation of $\mathbb{R}$ on a real Hilbert space $\mathsf{H}_{\mathbb{R}}$ and let $A$ denote the analytic generator of the complexification of $\{U_{t}\}_{t\in\mathbb{R}}$. Let $\mathsf{H}_{\mathbb{R}}^{ap}$ denote the closed real subspace generated by the eigenvectors of $A$ and let $\mathsf{H}_{\mathbb{R}}^{wm}={\mathsf{H}_{\mathbb{R}}^{ap}}^{\perp}$. The subspaces $\mathsf{H}_{\mathbb{R}}^{ap}$ and $\mathsf{H}_{\mathbb{R}}^{wm}$ are reducing under $A$ and every $\{U_{t}\}_{t\in\mathbb{R}}$, and they are respectively called the almost periodic and weakly mixing part of $\mathsf{H}_{\mathbb{R}}$. Similarly, ${U_{t}}_{|_{\mathsf{H}_{\mathbb{R}}^{ap}}}$ and ${U_{t}}_{|_{\mathsf{H}_{\mathbb{R}}^{wm}}}$ are called the almost periodic and weakly mixing part of $\{U_{t}\}_{t\in\mathbb{R}}$ respectively.
Note that whenever $\mathsf{H}_{\mathbb{R}}^{wm}\neq 0$, it must be infinite-dimensional. On the other hand, the almost periodic part decomposes as follows (see [Shl]):
$(\mathsf{H}_{\mathbb{R}}^{ap},{U_{t}}_{|_{\mathsf{H}_{\mathbb{R}}^{ap}}})=\left(\bigoplus_{j=1}^{N_{1}}(\mathbb{R},id)\right)\oplus\left(\bigoplus_{j=1}^{N_{2}}\mathsf{H}_{\mathbb{R}}^{k},U_{t}^{k}\right)$
for some $0\leq N_{1},N_{2}\leq\aleph_{0}$, where for all $1\leq k\leq N_{2}$ we have
$\mathsf{H}_{\mathbb{R}}^{k}=\mathbb{R}^{2},\;\;\;\;U_{t}^{k}=\begin{pmatrix}\cos(t\log\lambda_{k})&-\sin(t\log\lambda_{k})\\ \sin(t\log\lambda_{k})&\cos(t\log\lambda_{k})\end{pmatrix}$
for some $\lambda_{k}\in(0,1)$.
3. Technical analysis
This section contains the rigorous technical analysis, where we will prove a series of Lemmas and Propositions about various convergence results. The most crucial results are Propositions 3.8, 3.4 and 3.11. We recommend that readers first review the statements of these Propositions and then proceed directly to the proof of the main theorem in the next section to understand the reasons behind these calculations. The notations used in Proposition 3.8 are mentioned just before Lemma 3.6.
Lemma 3.1.
For any fixed non-zero vector $\xi\in\mathsf{H}_{U}$, $n\in\mathbb{N}$ and $1\leq m\leq n$, we have
(2) $c^{*}(\xi)^{m}c(\xi)^{n}=\frac{[n]_{q}!}{[n-m]_{q}!}\|\xi\|_{U}^{2m}c(\xi)^{n-m}+\sum_{k=0}^{m-1}\frac{[n]_{q}!}{[n-k]_{q}!}q^{n-k}\|\xi\|_{U}^{2k}c^{*}(\xi)^{m-k-1}c(\xi)^{n-k}c^{*}(\xi).$
Proof.
The proof is by induction on $m$. For $m=1$, use Lemma 2.3 to write
(3) $c^{*}(\xi)c(\xi)^{n}=[n]_{q}\|\xi\|_{U}^{2}c(\xi)^{n-1}+q^{n}c(\xi)^{n}c^{*}(\xi)$
which is exactly as in (2). Now assume the desired equation to be true for some $11\leq m\leq n-$ 1; then calculate
$c^{*}(\xi)^{m+1}c(\xi)^{n}$ $=c^{*}(\xi)\left(c^{*}(\xi)^{m}c(\xi)^{n}\right)$
$=\frac{[n]_{q}!}{[n-m]_{q}!}\|\xi\|_{U}^{2m}c^{*}(\xi)c(\xi)^{n-m}+\sum_{k=0}^{m-1}\frac{[n]_{q}!}{[n-k]_{q}!}\|\xi\|_{U}^{2k}q^{n-k}c^{*}(\xi)^{m-k}c(\xi)^{n-k}c^{*}(\xi)$
[…]
$=\frac{[n]_{q}!}{[n-m]_{q}!}\|\xi\|_{U}^{2m}\left([n-m]_{q}\|\xi\|_{U}^{2}c(\xi)^{n-m-1}+q^{n-m}c(\xi)^{n-m}c^{*}(\xi)\right)$
[…]
$=\frac{[n]_{q}!}{[n-m-1]_{q}!}\|\xi\|_{U}^{2m+2}c(\xi)^{n-(m+1)}+\sum_{k=0}^{m}\frac{[n]_{q}!}{[n-k]_{q}!}\|\xi\|_{U}^{2k}q^{n-k}c^{*}(\xi)^{m-k}c(\xi)^{n-k}c^{*}(\xi).$
The proof then ends by the induction argument. ∎
Corollary 3.2.
For vectors $\xi,\eta\in\mathsf{H}_{U}$ with $\langle\xi,\eta\rangle_{U}=0$, $n\in\mathbb{N}$, $0\leq m\leq n$ and $\ell\geq 1$, we have
$c^{*}(\xi)^{m}c(\xi)^{n}c(\eta)^{\ell}=\frac{[n]_{q}!}{[n-m]_{q}!}\|\xi\|_{U}^{2m}c(\xi)^{n-m}c(\eta)^{\ell}+q^{\ell}X$
where $X$ is an operator bounded by $A_{q}^{\prime}(1-q)^{-\frac{(m+n+\ell)}{2}}\|\xi\|_{U}^{m+n}\|\eta\|_{U}^{\ell}$ for some constant $A_{q}^{\prime}$ depending only on $q$.
Proof.
If $m=0$ then set $X=0$, and if $m\geq 1$ then set
$X=\sum_{k=0}^{m-1}\frac{[n]_{q}!}{[n-k]_{q}!}q^{n-k}\|\xi\|_{U}^{2k}c^{*}(\xi)^{m-k-1}c(\xi)^{n-k}c(\eta)^{\ell}c^{*}(\xi).$
Obtain the required expression by multiplying from the right in eq. (2) of Lemma 3.1 by $c(\eta)^{\ell}$ and then by noting via $q$-commutation relation that $c^{*}(\xi)c(\eta)^{\ell}=q^{\ell}c(\eta)^{\ell}c^{*}(\xi)$. Further, we estimate the norm of $X$ using Lemma 2.4 as follows:
$\|X\|$ $\leq C_{q}^{4}\sum_{k=0}^{m-1}\frac{[n]_{q}!}{[n-k]_{q}!}|q|^{n-k}\|\xi\|_{U}^{2k}\|\xi\|_{U}^{(m-k-1)+(n-k)+1}\|\eta\|_{U}^{\ell}(1-q)^{-\frac{1}{2}(m-k-1)-\frac{1}{2}(n-k)-\frac{\ell}{2}-\frac{1}{2}}$
$=C_{q}^{4}\sum_{k=0}^{m-1}\frac{d_{n}(1-q)^{-n}}{d_{n-k}(1-q)^{-n+k}}|q|^{n-k}\|\xi\|_{U}^{m+n}\|\eta\|_{U}^{\ell}(1-q)^{-\frac{(m+n+\ell)}{2}+k}$
$\leq C_{q}^{6}\|\xi\|_{U}^{m+n}\|\eta\|_{U}^{\ell}(1-q)^{-\frac{(m+n+\ell)}{2}}\sum_{k=0}^{m-1}|q|^{n-k}\quad\quad\quad\quad(\text{as }d_{n},\frac{1}{d_{n-k}}\leq C_{q})$
$=\|\xi\|_{U}^{m+n}\|\eta\|_{U}^{\ell}(1-q)^{-\frac{(m+n+\ell)}{2}}\frac{C_{q}^{6}}{1-|q|}$
which gives the required estimate by setting $A_{q}^{\prime}=\frac{C_{q}^{6}}{1-|q|}$. ∎
Corollary 3.3.
For vectors $\xi,\eta\in\mathsf{H}_{U}$ with $\langle\xi,\eta\rangle_{U}=0$, and $m>n\geq 1$, $\ell\geq 1$, we have
$c^{*}(\xi)^{m}c(\xi)^{n}c(\eta)^{\ell}=q^{\ell}Y$
where $Y$ is an operator bounded by $A_{q}(1-q)^{-\frac{(m+n+\ell)}{2}}\|\xi\|_{U}^{m+n}\|\eta\|_{U}^{\ell}$ for some constant $A_{q}$ depending only on $q$.
Proof.
Using Corollary 3.2, write
$c^{*}(\xi)^{n}c(\xi)^{n}c(\eta)^{\ell}=[n]_{q}!\|\xi\|_{U}^{2n}c(\eta)^{\ell}+q^{\ell}X$
where $X$ is an operator bounded by $\frac{C_{q}^{6}}{1-|q|}(1-q)^{-n-\frac{\ell}{2}}\|\xi\|_{U}^{2n}\|\eta\|_{U}^{\ell}$. By multiplying both sides from the left by $c^{*}(\xi)^{m-n}$ and using the $q$-commutation relation $c^{*}(\xi)^{m-n}c(\eta)^{\ell}=q^{(m-n)\ell}c(\eta)^{\ell}c^{*}(\xi)^{(m-n)}$ since $\langle\xi,\eta\rangle_{U}=0$, we get
$c^{*}(\xi)^{m}c(\xi)^{n}c(\eta)^{\ell}=q^{(m-n)\ell}[n]_{q}!\|\xi\|_{U}^{2n}c(\eta)^{\ell}c^{*}(\xi)^{m-n}+q^{\ell}c^{*}(\xi)^{m-n}X=q^{\ell}Y$
where $Y=q^{(m-n-1)\ell}[n]_{q}!\|\xi\|_{U}^{2n}c(\eta)^{\ell}c^{*}(\xi)^{m-n}+c^{*}(\xi)^{m-n}X$. It is an easy verify as before using Lemma 2.4 that
$\|Y\|$ $\leq|q|^{(m-n-1)\ell}d_{n}C_{q}^{2}\|\xi\|_{U}^{m+n}\|\eta\|_{U}^{\ell}(1-q)^{-\frac{(m+n+\ell)}{2}}+\frac{C_{q}^{6}}{1-|q|}\|\xi\|_{U}^{m+n}\|\eta\|_{U}^{\ell}C_{q}(1-q)^{-\frac{(m+n+\ell)}{2}}$
$\leq A_{q}\|\xi\|_{U}^{m+n}\|\eta\|_{U}^{\ell}(1-q)^{-\frac{(m+n+\ell)}{2}}$
where $A_{q}=|q|^{(m-n-1)\ell}C_{q}^{3}+\frac{C_{q}^{7}}{1-|q|}$. ∎
Proposition 3.4.
For any non-zero vectors $\xi,\eta\in\mathsf{H}_{U}$ with $\langle\xi,\eta\rangle_{U}=0$ and a uniformly bounded sequence $\{T_{k}\}_{k\geq 1}$ of operators, the sequences
  1. (1)
    $\frac{1}{n}\sum_{k=1}^{n}(1-q)^{k}\|\xi\|_{U}^{-k}\|\eta\|_{U}^{-k}c(\xi)^{k}c(\eta)^{k}T_{k}$
  2. (2)
    $\frac{1}{n}\sum_{k=1}^{n}(1-q)^{k}\|\xi\|_{U}^{-k}\|\eta\|_{U}^{-k}r(\xi)^{k}r(\eta)^{k}T_{k}$
converge to 00 in norm as $n\to\infty$.
Proof.
We shall prove the convergence of the first sequence, which we call $X_{n}$. The proof for the second sequence follows similarly. By replacing $\xi$ and $\eta$ with $\frac{\xi}{\|\xi\|_{U}}$ and $\frac{\eta}{\|\eta\|_{U}}$ respectively, we may and will assume without loss of generality that $\xi$ and $\eta$ are unit vectors. Also assume that $\sup_{k\geq 1}\|T_{k}\|=1$. Now calculate
$X_{n}^{*}X_{n}$ $=\frac{1}{n^{2}}\sum_{\ell=1}^{n}\sum_{k=1}^{n}(1-q)^{\ell+k}T_{\ell}^{*}c^{*}(\eta)^{\ell}c^{*}(\xi)^{\ell}c(\xi)^{k}c(\eta)^{k}T_{k}$
$=X^{n}_{\ell=k}+X^{n}_{k<\ell}+X^{n}_{k>\ell}$
where we have divided the sum over three cases: $\ell=k,k<\ell$ and $k>\ell$ i.e.
$X_{\ell=k}^{n}=\frac{1}{n^{2}}\sum_{k=1}^{n}(1-q)^{2k}T_{k}^{*}c^{*}(\eta)^{k}c^{*}(\xi)^{k}c(\xi)^{k}c(\eta)^{k}T_{k}$
$X_{k<\ell}^{n}=\frac{1}{n^{2}}\sum_{1\leq k<\ell\leq n}(1-q)^{\ell+k}T_{\ell}^{*}c^{*}(\eta)^{\ell}c^{*}(\xi)^{\ell}c(\xi)^{k}c(\eta)^{k}T_{k}$
$X_{\ell<k}^{n}=\frac{1}{n^{2}}\sum_{n\geq k>\ell\geq 1}(1-q)^{\ell+k}T_{\ell}^{*}c^{*}(\eta)^{\ell}c^{*}(\xi)^{\ell}c(\xi)^{k}c(\eta)^{k}T_{k}.$
We now estimate the norm of the three sums separately. Firstly, over $k=\ell$, we estimate the norm using Lemma 2.4 as follows:
$\|X_{\ell=k}^{n}\|\leq\frac{1}{n^{2}}\sum_{k=1}^{n}C_{q}^{4}=\frac{C_{q}^{4}}{n}.$
We next estimate the norm of $X_{k<\ell}^{n}$: for $k<\ell$, we use Corollary 3.3 to write $c^{*}(\xi)^{\ell}c(\xi)^{k}c(\eta)^{k}=q^{k}X_{k,\ell}$ where $X_{k,\ell}$ is an operator bounded by $A_{q}(1-q)^{-\frac{\ell}{2}-k}$. Therefore, we have
$\|X_{k<\ell}^{n}\|$ $=\frac{1}{n^{2}}\left\|\sum_{1\leq k<\ell\leq n}(1-q)^{\ell+k}c^{*}(\eta)^{\ell}T_{\ell}^{*}\left(q^{k}X_{k,\ell}\right)T_{k}\right\|$
$\leq\frac{C_{q}A_{q}}{n^{2}}\sum_{1\leq k<\ell\leq n}|q|^{k}(1-q)^{\ell+k}(1-q)^{-\frac{\ell}{2}-k}(1-q)^{-\frac{\ell}{2}}$
$\leq\frac{C_{q}A_{q}}{n^{2}}\sum_{k,\ell=1}^{n}|q|^{k}$
$\leq\frac{C_{q}A_{q}}{n}\sum_{k=1}^{\infty}|q|^{k}$
$\leq\frac{1}{n}\cdot\frac{C_{q}A_{q}}{1-|q|}.$
Similarly, by symmetry $X_{k>\ell}^{n}=(X_{k<\ell}^{n})^{*}$, $\|X^{n}_{k>\ell}\|$ has the same estimate $\|X_{k<\ell}^{n}\|$. Thus we obtain that
$\|X_{n}^{*}X_{n}\|\leq\frac{1}{n}\left({C_{q}^{4}}+\frac{2C_{q}A_{q}}{1-|q|}\right)$
which proves that $X_{n}\to 0$ in norm as $n\to\infty$. ∎
Remark 3.5.
It is very important that we put the operator $T_{k}$ on the right side in Proposition 3.4 and not on the left. In fact, the operator $\tilde{S}_{n}=\frac{1}{n}\sum_{k=0}^{n}(1-q)^{k}\tilde{T}_{k}c(\xi)^{k}c(\eta)^{k}$ need not converge to 00 for fixed unit vectors $\xi,\eta\in\mathsf{H}_{U}$ with $\langle\xi,\eta\rangle_{U}=0$ and bounded operators $\tilde{T}_{k}$. See for example Proposition 3.8 below.
We now restrict our attention to a two-dimensional invariant subspace of a given representation $\{U_{t}\}_{t\in\mathbb{R}}$. Let $\mathsf{K}_{\mathbb{R}}=\mathbb{R}^{2}$ be an invariant subspace of $\{U_{t}\}_{t\in\mathbb{R}}$ such that ${U_{t}}_{|_{\mathsf{K}_{\mathbb{R}}}}$ is ergodic i.e. has no fixed nonzero vector. Then, as discussed in Section 2, there is a scalar $\lambda\in(0,1)$ such that
${U_{t}}_{|_{\mathsf{K}_{\mathbb{R}}}}=\begin{pmatrix}\cos(t\log\lambda)&-\sin(t\log\lambda)\\ \sin(t\log\lambda)&\cos(t\log\lambda)\end{pmatrix}.$
Set
$e=\frac{({\lambda+1})^{\frac{1}{2}}}{2\lambda^{\frac{1}{4}}}\begin{pmatrix}1\\ i\end{pmatrix}\in\mathsf{K}_{\mathbb{C}}=\mathbb{C}^{2}.$
If $A$ denotes the analytic generator, then it is immediate to check that
$Ae=\lambda^{-1}e,\;\;\;A\bar{e}=\lambda\bar{e},\;\;\;\|e\|_{U}=\lambda^{-\frac{1}{4}},\;\;\;\|\bar{e}\|_{U}=\lambda^{\frac{1}{4}}\;\;\;\mbox{ and }\langle e,\bar{e}\rangle_{U}=0.$
Moreover, the (right) conjugate of $e$ and $\bar{e}$ with respect to the decomposition in $\mathsf{H}_{\mathbb{R}}^{\prime}+i\mathsf{H}_{\mathbb{R}}^{\prime}$ is given as follows:
$\bar{e}^{r}=\lambda^{-1}\bar{e}\;\;\;\mbox{ and }\;\;\;\overline{(\bar{e})}^{r}=\lambda e.$
Indeed it is straightforward to check that $(1+A^{-1})^{-1}(e+\lambda^{-1}\bar{e})=\frac{1}{1+\lambda}(e+\bar{e})\in\mathsf{H}_{\mathbb{R}}$ and hence $\langle e+\lambda^{-1}\bar{e},\xi\rangle_{U}\in\mathbb{R}$ for all $\xi\in\mathsf{H}_{\mathbb{R}}$, and similarly check that $\langle e-\lambda^{-1}\bar{e},\xi\rangle_{U}\in i\mathsf{H}_{\mathbb{R}}$ for all $\xi\in\mathsf{H}_{\mathbb{R}}$; this will prove that $e+\lambda^{-1}\bar{e}\in\mathsf{H}_{\mathbb{R}}^{\prime}$ and $e-\lambda^{-1}\bar{e}\in i\mathsf{H}_{\mathbb{R}}^{\prime}$ and hence $\bar{e}^{r}=\lambda^{-1}\bar{e}$.
We fix such $\lambda,e$ and $\bar{e}$ for the rest of this section. The notations used in what follows have been explained in Section 2. Our next main result is Proposition 3.8, but first we need two crucial lemmas. The following lemma is borrowed from [BMRW].
Lemma 3.6.
The sequences $\{(1-q)^{n}W(\bar{e}^{n}e^{n})\}_{n\geq 1}$ and $\{(1-q)^{n}W_{r}(\bar{e}^{n}e^{n})\}_{n\geq 1}$ are bounded.
Proof.
We briefly sketch the proof from [BMRW, Lemma 3.8] in order to set some notations and record some useful observations. Note that both the sequences have the same norm; indeed we have $W_{r}(\bar{e}^{n}e^{n})=J_{\varphi}W(\bar{e}^{n}e^{n})J_{\varphi}$. We consider the first sequence which by Wick formula is given by
(4) $W(\bar{e}^{n}e^{n})=\sum_{k,\ell=0}^{n}q_{k,\ell}^{(n)}c(\bar{e})^{k}c(e)^{\ell}c^{*}(e)^{n-k}c^{*}(\bar{e})^{n-\ell}$
where $q_{k,\ell}^{(n)}=\sum_{J_{1},J_{2}}q^{c(J,J^{c})}$ with the sum running over all possible subsets $J_{1}\subseteq\{1,\ldots,n\}$ of cardinality $k$ and $J_{2}\subseteq\{n+1,\ldots,2n\}$ of cardinality $\ell$ such that $J=J_{1}\cup J_{2}$. Here $c(J,J^{c})$ denotes the number of crossings of the partition $J\cup J^{c}$ of $\{1,2,\ldots,2n\}$ (see [BMRW, Lemma 3.8] for more explanation of the terms $q_{k,\ell}^{(n)}$). Then we have $|q_{k,\ell}^{(n)}|\leq C_{q}^{2}|q|^{(n-k)\ell}$ and
$(1-q)^{n}\sum_{k,\ell=0}^{n}|q_{k,\ell}^{(n)}|\;\|c(\bar{e})^{k}\|\;\|c(e)^{\ell}\|\;\|c^{*}(e)^{n-k}\|\;\|c^{*}(\bar{e})^{n-\ell}\|\leq C_{q}^{6}\sum_{k,\ell=0}^{n}|q|^{(n-k)\ell}\lambda^{\frac{k-\ell}{2}}.$
The important observation is that $\sum_{k,\ell=0}^{n}|q|^{(n-k)\ell}\lambda^{\frac{k-\ell}{2}}\leq\frac{1}{(1-|q|)(1-\lambda^{\frac{1}{2}})}+\frac{(n+1)|q|^{n}\lambda^{-\frac{1}{2}}}{1-|q|^{n}\lambda^{-\frac{1}{2}}}$ as $n\to\infty$, which means that there is a constant $B(q,\lambda)$ depending on $q$ and $\lambda$ such that
$\sum_{k,\ell=0}^{n}|q|^{(n-k)\ell}\lambda^{\frac{k-\ell}{2}}\leq B(q,\lambda)$
for all $n\geq 1$. This observation will be useful for us later as well (also see Lemma 3.7 below). ∎
Lemma 3.7.
For any $q\in(-1,1)$ and $\lambda\in(0,1)$, the sequence
$s_{n}=\frac{1}{n}\sum_{m=1}^{n}\sum_{\ell=1}^{m}\sum_{k=0}^{m-1}q^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}$
converges to 00 as $n\to\infty$.
Proof.
Assume that $q\geq 0$. Write $s_{n}$ as follows:
$s_{n}$ $=\frac{1}{n}\sum_{m=1}^{n}\sum_{0\leq k<\ell\leq m}q^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}+\frac{1}{n}\sum_{m=1}^{n}\sum_{1\leq\ell\leq k\leq m-1}q^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}$
(5) $=\frac{1}{n}\sum_{m=1}^{n}\sum_{j=-m}^{-1}\sum_{k=0}^{m+j}q^{(m-k)(k-j)}\lambda^{\frac{j}{2}}+\frac{1}{n}\sum_{m=1}^{n}\sum_{j=0}^{m-1}\sum_{k=j+1}^{m-1}q^{(m-k)(k-j)}\lambda^{\frac{j}{2}}$
where we get the last equality by putting $j=k-\ell$. Now we estimate the first quantity in (5), where we follow the arguments as in the proof of [BMRW, Lemma 3.9]:
$\frac{1}{n}\sum_{m=1}^{n}\sum_{j=-m}^{-1}\sum_{k=0}^{m+j}q^{(m-k)(k-j)}\lambda^{\frac{j}{2}}$ $=\frac{1}{n}\sum_{m=1}^{n}\sum_{j=1}^{m}\lambda^{-\frac{j}{2}}\sum_{k=0}^{m-j}q^{(m-k)(k+j)}\leq\frac{1}{n}\sum_{m=1}^{n}\sum_{j=1}^{m}\lambda^{-\frac{j}{2}}(m-j+1)q^{mj}.$
Choose $n_{0}$ large enough such that $q^{n_{0}}\lambda^{-\frac{1}{2}}\leq\frac{1}{2}$; hence for $m\geq n_{0}$ we have $q^{m}\lambda^{-\frac{1}{2}}\leq\frac{1}{2}q^{m-n_{0}}$ so that the above sum further yields
$\frac{1}{n}\sum_{m=1}^{n}\sum_{j=-m}^{-1}\sum_{k=0}^{m+j}q^{(m-k)(k-j)}\lambda^{\frac{j}{2}}$ $\leq\frac{1}{n}\sum_{m=1}^{n_{0}}\sum_{j=1}^{m}\lambda^{-\frac{j}{2}}(m-j+1)q^{mj}+\frac{1}{n}\sum_{m=n_{0}+1}^{n}\sum_{j=1}^{m}(m+1)\left(q^{m}\lambda^{-\frac{1}{2}}\right)^{j}$
$\leq\frac{C(q,\lambda,n_{0})}{n}+\frac{1}{n}\sum_{m=n_{0}+1}^{n}(m+1)\sum_{j=1}^{m}\left(\frac{q^{m-n_{0}}}{2}\right)^{j}$
$\leq\frac{C(q,\lambda,n_{0})}{n}+\frac{1}{n}\sum_{m=1}^{n}(m+1)\frac{q^{m-n_{0}}}{2-q^{m-n_{0}}}$
$\to 0\mbox{ as }n\to\infty,$
where $C(q,\lambda,n_{0})$ is a constant independent from $n$, and where we have also used the fact that the sequence $\left\{(m+1)\frac{q^{m}}{2q^{n_{0}}-q^{m}}\right\}_{m\geq 1}$ converges to 00 as $m\to\infty$ and hence its Cesáro sum also converges to 00. On the other hand, the second quantity in (5) is estimated as follows:
$\frac{1}{n}\sum_{m=1}^{n}\sum_{j=0}^{m-1}\sum_{k=j+1}^{m-1}q^{(m-k)(k-j)}\lambda^{\frac{j}{2}}$ $=\frac{1}{n}\sum_{j=0}^{n-1}\lambda^{\frac{j}{2}}\sum_{k=j+1}^{n-1}\sum_{m=k+1}^{n}q^{(m-k)(k-j)}$
$=\frac{1}{n}\sum_{j=0}^{n-1}\lambda^{\frac{j}{2}}\sum_{k=j+1}^{n-1}q^{-k(k-j)}\sum_{m=k+1}^{n}q^{m(k-j)}$
$=\frac{1}{n}\sum_{j=0}^{n-1}\lambda^{\frac{j}{2}}\sum_{k=j+1}^{n-1}q^{-k(k-j)}\left(\frac{q^{(k+1)(k-j)}-q^{(n+1)(k-j)}}{1-q^{k-j}}\right)$
$\leq\frac{1}{n}\sum_{j=0}^{n-1}\lambda^{\frac{j}{2}}\sum_{k=j+1}^{n-1}\frac{q^{k-j}}{1-q^{k-j}}$
$\leq\frac{1}{1-\lambda^{\frac{1}{2}}}\left(\frac{1}{n}\sum_{k=1}^{n}\frac{q^{k}}{1-q^{k}}\right)$
$\to 0\mbox{ as }n\to\infty$
as the sequence $\{\frac{q^{k}}{1-q^{k}}\}_{k\geq 1}$ converges to 00 as $k\to\infty$, and so does its Cesáro sum. Thus we have proved that $s_{n}\to 0$ as $n\to\infty$. ∎
Proposition 3.8.
Let $e,\bar{e}$ and $\lambda$ be as above. The sequence
$S_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}W(\bar{e}^{m}e^{m})c(\bar{e})^{m}c(e)^{m}$
is convergent in norm to the operator $S$ given by
(6) $S=\sum_{k=0}^{\infty}\frac{d_{\infty}}{d_{k}}(1-q)^{k}\;\lambda^{\frac{k}{2}}c(\bar{e})^{k}Tc(e)^{k}$
where $T$ is the norm limit of the operator $(1-q)^{n}\lambda^{\frac{n}{2}}c^{*}(e)^{n}c(e)^{n}$ which is positive and invertible. Moreover, $S$ is invertible if $\lambda<\left(1+C_{q}^{4}\right)^{-2}$.
Proof.
By [BMRW, Lemma 3.4], the operator $T$ exists and is both positive and invertible. First check that the series defined in (6) above is norm convergent so that $S$ is a well-defined bounded operator. Indeed if we set $Z_{k}=\frac{d_{\infty}}{d_{k}}(1-q)^{k}\;\lambda^{\frac{k}{2}}c(\bar{e})^{k}Tc(e)^{k}$ for $k\geq 0$ then by using Lemma 2.5 we have
$\sum_{k=0}^{\infty}\|Z_{k}\|$ $\leq\sum_{k=0}^{\infty}\;\frac{d_{\infty}}{d_{k}}(1-q)^{k}\;\lambda^{\frac{k}{2}}\|c(\bar{e})^{k}\|\;\|T\|\;\|c(e)^{k}\|$
$\leq d_{\infty}C_{q}\;\|T\|\sum_{k=0}^{\infty}(1-q)^{k}\;\lambda^{\frac{k}{2}}\left(C_{q}(1-q)^{-\frac{k}{2}}\lambda^{\frac{k}{4}}\right)\left(C_{q}(1-q)^{-\frac{k}{2}}\lambda^{-\frac{k}{4}}\right)$
$=d_{\infty}C_{q}^{3}\;\|T\|\sum_{k=0}^{\infty}\lambda^{\frac{k}{2}}$
<<$\infty$.< ∞ .
To prove the convergence of $S_{n}$, write using (4)
$S_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{k,\ell=0}^{m}q_{k,\ell}^{(m)}c(\bar{e})^{k}c(e)^{\ell}c^{*}(e)^{m-k}c^{*}(\bar{e})^{m-\ell}c(\bar{e})^{m}c(e)^{m}$
where $q_{k,\ell}^{(m)}$ are the scalars as in the proof of Lemma 3.6 which satisfy $|q^{(m)}_{k,\ell}|\leq C_{q}^{2}|q|^{(m-k)\ell}$. Let us understand the terms $c^{*}(\bar{e})^{m-\ell}c(\bar{e})^{m}c(e)^{m}$ in the sum above. Use Corollary 3.2 to write
$c^{*}(\bar{e})^{m-\ell}c(\bar{e})^{m}c(e)^{m}=\frac{[m]_{q}!}{[\ell]_{q}!}\lambda^{\frac{m-\ell}{2}}c(\bar{e})^{\ell}c(e)^{m}+q^{m}X_{m,\ell}$
where $X_{m,\ell}$ is an operator bounded by $(1-q)^{\frac{(-3m+\ell)}{2}}\lambda^{\frac{m-\ell}{4}}A_{q}$ for the constant $A_{q}:=\frac{C_{q}^{6}}{1-|q|}$ depending only on $q$. In particular, we get
$S_{n}$ $=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{k,\ell=0}^{m}q_{k,\ell}^{(m)}\left(\frac{[m]_{q}!}{[\ell]_{q}!}\lambda^{\frac{m-\ell}{2}}\right)c(\bar{e})^{k}c(e)^{\ell}c^{*}(e)^{m-k}c(\bar{e})^{\ell}c(e)^{m}$
$\quad\quad\quad\quad+\frac{1}{n}\sum_{m=1}^{n}q^{m}(1-q)^{2m}\sum_{k,\ell=0}^{m}q_{k,\ell}^{(m)}c(\bar{e})^{k}c(e)^{\ell}c^{*}(e)^{m-k}X_{m,\ell}.$
We now calculate the upper bound for the last sum above:
$\frac{1}{n}\sum_{m=1}^{n}|q|^{m}(1-q)^{2m}$ $\left\|\sum_{k,\ell=0}^{m}q_{k,\ell}^{(m)}c(\bar{e})^{k}c(e)^{\ell}c^{*}(e)^{m-k}X_{m,\ell}\right\|$
$\leq\frac{1}{n}\sum_{m=1}^{n}|q|^{m}(1-q)^{2m}C_{q}^{5}A_{q}\sum_{k,\ell=0}^{m}|q|^{(m-k)\ell}\lambda^{\frac{k-\ell-(m-k)}{4}}(1-q)^{-\frac{(k+\ell+m-k)}{2}}(1-q)^{\frac{(-3m+\ell)}{2}}\lambda^{\frac{m-\ell}{2}}$
$=\frac{1}{n}C_{q}^{5}A_{q}\sum_{m=1}^{n}|q|^{m}\sum_{k,\ell=0}^{m}|q|^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}$
$\leq\frac{1}{n}\frac{C_{q}^{5}A_{q}B(q,\lambda)}{1-|q|}$
where $B(q,\lambda)$ is an upper bound for the sum $\sum_{k,\ell=0}^{m}|q|^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}$ such that $B(q,\lambda)$ depends only on $\lambda$ and $q$, and is independent of $m$ (see the proof of Lemma 3.6). Thus we have
$S_{n}$ $=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{k,\ell=0}^{m}q_{k,\ell}^{(m)}\left(\frac{[m]_{q}!}{[\ell]_{q}!}\lambda^{\frac{m-\ell}{2}}\right)c(\bar{e})^{k}c(e)^{\ell}c^{*}(e)^{m-k}c(\bar{e})^{\ell}c(e)^{m}+O\left(\frac{1}{n}\right)$
$=X_{n}+Y_{n}+Z_{n}+O\left(\frac{1}{n}\right)$
where the above sum is divided over $\ell=0,k=m$ and the rest i.e. we write
$X_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{k=0}^{m}q_{k,0}^{(m)}[m]_{q}!\lambda^{\frac{m}{2}}c(\bar{e})^{k}c^{*}(e)^{m-k}c(e)^{m}$
$Y_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{\ell=0}^{m}q_{m,\ell}^{(m)}\frac{[m]_{q}!}{[\ell]_{q}!}\lambda^{\frac{m-\ell}{2}}c(\bar{e})^{m}c(e)^{\ell}c(\bar{e})^{\ell}c(e)^{m}$
$Z_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{\ell=1}^{m}\sum_{k=0}^{m-1}q_{k,\ell}^{(m)}\left(\frac{[m]_{q}!}{[\ell]_{q}!}\lambda^{\frac{m-\ell}{2}}\right)c(\bar{e})^{k}c(e)^{\ell}c^{*}(e)^{m-k}c(\bar{e})^{\ell}c(e)^{m}.$
But the norm of $Z_{n}$ can be estimated using Lemma 2.4 as follows:
$\|Z_{n}\|$ $\leq\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{\ell=1}^{m}\sum_{k=0}^{m-1}C_{q}^{2}|q|^{(m-k)\ell}\left(\frac{d_{m}}{d_{\ell}}\lambda^{\frac{m-\ell}{2}}\right)C_{q}^{5}(1-q)^{-2m}\lambda^{\frac{k-m}{2}}$
$\leq\frac{C_{q}^{9}}{n}\sum_{m=1}^{n}\sum_{\ell=1}^{n}\sum_{k=0}^{m-1}|q|^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}$
$=C_{q}^{9}s_{n}$
where $s_{n}$ is the sequence as in Lemma 3.7; hence $Z_{n}\to 0$ as $n\to\infty$ by the same Lemma. We now show that $Y_{n}$ converges to 0 as $n\to\infty$. We follow the same trick as in the proof of Proposition 3.4. Calculate
$Y_{n}^{*}Y_{n}$ $=\frac{1}{n^{2}}\sum_{m,m^{\prime}=1}^{n}(1-q)^{2m+2m^{\prime}}$
$\cdot\sum_{\ell=0}^{m}\sum_{\ell^{\prime}=0}^{m^{\prime}}q_{m,\ell}^{(m)}q_{m^{\prime},\ell^{\prime}}^{(m^{\prime})}\frac{[m]_{q}![m^{\prime}]_{q}}{[\ell]_{q}![\ell^{\prime}]_{q}!}\lambda^{\frac{m+m^{\prime}-\ell-\ell^{\prime}}{2}}c^{*}(e)^{m}c^{*}(\bar{e})^{\ell}c^{*}(e)^{\ell}c^{*}(\bar{e})^{m}c(\bar{e})^{m^{\prime}}c(e)^{\ell^{\prime}}c(\bar{e})^{\ell^{\prime}}c(e)^{m^{\prime}}$
$=W^{n}_{m=m^{\prime}}+W^{n}_{m>m^{\prime}}+W^{n}_{m<m^{\prime}}$
where we divide the sum over three cases: $m=m^{\prime},m<m^{\prime}$ and $m>m^{\prime}$. Over $m=m^{\prime}$, we again use some similar estimates to calculate the following:
$\|W_{m=m^{\prime}}^{n}\|$ $=\frac{1}{n^{2}}\left\|\sum_{m=1}^{n}(1-q)^{4m}\sum_{\ell,\ell^{\prime}=0}^{m}q_{m,\ell}^{(m)}q_{m,\ell^{\prime}}^{(m)}\frac{([m]_{q}!)^{2}}{[\ell]_{q}![\ell^{\prime}]_{q}!}\lambda^{\frac{2m-\ell-\ell^{\prime}}{2}}c^{*}(e)^{m}c^{*}(\bar{e})^{\ell}c^{*}(e)^{\ell}c^{*}(\bar{e})^{m}c(\bar{e})^{m}c(e)^{\ell^{\prime}}c(\bar{e})^{\ell^{\prime}}c(e)^{m}\right\|$
$\leq\frac{C_{q}^{12}}{n^{2}}\sum_{m=1}^{n}(1-q)^{4m}\sum_{\ell,\ell^{\prime}=0}^{m}\frac{d_{m}^{2}(1-q)^{-2m}}{d_{\ell}d_{\ell^{\prime}}(1-q)^{-\ell-\ell^{\prime}}}\lambda^{\frac{2m-\ell-\ell^{\prime}}{2}}(1-q)^{-2m-\ell-\ell^{\prime}}\quad\quad\quad\quad(\text{since }q_{m,\ell}^{(m)}\leq C_{q}^{2})$
$\leq\frac{C_{q}^{16}}{n^{2}}\sum_{m=1}^{n}\sum_{\ell,\ell^{\prime}=0}^{m}\lambda^{\frac{2m-\ell-\ell^{\prime}}{2}}$
$\leq\frac{1}{n}\frac{C_{q}^{16}}{(1-\sqrt{\lambda})^{2}}.$
Next consider the sum over $m>m^{\prime}$ (which has the same norm as the part $m<m^{\prime}$ by considering the adjoint and the symmetric nature of the sum). For $m>m^{\prime}$, use Corollary 3.3 to write
$c^{*}(\bar{e})^{m}c(\bar{e})^{m^{\prime}}c(e)^{\ell^{\prime}}=q^{\ell^{\prime}}X_{m,m^{\prime},\ell^{\prime}}$
where $X_{m,m^{\prime},\ell^{\prime}}$ is an operator bounded by $A_{q}(1-q)^{-\frac{(m+m^{\prime}+\ell^{\prime})}{2}}\lambda^{\frac{m+m^{\prime}-\ell^{\prime}}{4}}$ for some constant $A_{q}$. Therefore, the sum over $m>m^{\prime}$ is estimated as
$\|W^{n}_{m>m^{\prime}}\|$
$=\frac{1}{n^{2}}\left\|\sum_{n\geq m>m^{\prime}\geq 1}^{n}(1-q)^{2m+2m^{\prime}}\sum_{\ell=0}^{m}\sum_{\ell^{\prime}=0}^{m^{\prime}}q_{m,\ell}^{(m)}q_{\ell^{\prime},m^{\prime}}^{(m^{\prime})}\frac{[m]_{q}![m^{\prime}]_{q}}{[\ell]_{q}![\ell^{\prime}]_{q}!}\lambda^{\frac{m+m^{\prime}-\ell-\ell^{\prime}}{2}}c^{*}(e)^{m}c^{*}(\bar{e})^{\ell}c^{*}(e)^{\ell}\left(q^{\ell^{\prime}}X_{m,m^{\prime},\ell^{\prime}}\right)c(\bar{e})^{\ell^{\prime}}c(e)^{m^{\prime}}\right\|$
$\leq\frac{C_{q}^{9}A_{q}}{n^{2}}\sum_{n\geq m>m^{\prime}\geq 1}(1-q)^{2m+2m^{\prime}}\sum_{\ell=0}^{m}\sum_{\ell^{\prime}=0}^{m^{\prime}}|q|^{\ell^{\prime}}\frac{d_{m}d_{m^{\prime}}(1-q)^{-m-m^{\prime}}}{d_{\ell}d_{\ell^{\prime}}(1-q)^{-\ell-\ell^{\prime}}}\lambda^{\frac{m+m^{\prime}-\ell-\ell^{\prime}}{2}}(1-q)^{-m-m^{\prime}-\ell-\ell^{\prime}}\quad(\text{as }q_{m,\ell}^{(m)}\leq C_{q}^{2})$
$\leq\frac{C_{q}^{13}A_{q}}{n^{2}}\sum_{m,m^{\prime}=1}^{n}\sum_{\ell=0}^{m}\sum_{\ell^{\prime}=0}^{m^{\prime}}|q|^{\ell^{\prime}}\lambda^{\frac{m+m^{\prime}-\ell-\ell^{\prime}}{2}}$
$\leq\frac{C_{q}^{13}A_{q}}{n^{2}}\frac{n}{1-\sqrt{\lambda}}\sum_{m^{\prime}=1}^{n}\sum_{\ell^{\prime}=0}^{m^{\prime}}|q|^{\ell^{\prime}}\lambda^{\frac{m^{\prime}-\ell^{\prime}}{2}}$
$\leq\frac{C_{q}^{13}A_{q}}{n(1-\sqrt{\lambda})}\sum_{m^{\prime}=1}^{n}\left(\sum_{\ell^{\prime}=0}^{\frac{m}{2}}|q|^{\ell^{\prime}}\lambda^{\frac{m^{\prime}-\ell^{\prime}}{2}}+\sum_{\ell^{\prime}=\frac{m^{\prime}}{2}}^{m}|q|^{\ell^{\prime}}\lambda^{\frac{m^{\prime}-\ell^{\prime}}{2}}\right)$
$\leq\frac{C_{q}^{13}A_{q}}{n(1-\sqrt{\lambda})}\left(\sum_{m^{\prime}=1}^{n}\lambda^{\frac{m^{\prime}}{2}}\sum_{\ell^{\prime}=0}^{\frac{m}{2}}|q|^{\ell^{\prime}}+\sum_{m^{\prime}=1}^{n}|q|^{\frac{m^{\prime}}{2}}\sum_{\ell^{\prime}=\frac{m^{\prime}}{2}}^{m}\lambda^{\frac{m^{\prime}-\ell^{\prime}}{2}}\right)$
$\leq\frac{C_{q}^{13}A_{q}}{n(1-\sqrt{\lambda})}\left(\frac{1}{(1-\sqrt{\lambda})(1-|q|)}+\frac{1}{(1-\sqrt{|q|})(1-\sqrt{\lambda})}\right)$
$\leq\frac{1}{n}\frac{2C_{q}^{13}A_{q}}{(1-\sqrt{\lambda})^{2}(1-\sqrt{|q|})}.$
A similar upper bound can be given for $\|W_{m<m^{\prime}}^{n}\|$. Thus we have shown that
$\|Y_{n}^{*}Y_{n}\|\leq\frac{1}{n}\left(\frac{C_{q}^{16}}{(1-\sqrt{\lambda})^{2}}+\frac{4C_{q}^{13}A_{q}}{(1-\sqrt{\lambda})^{2}(1-\sqrt{|q|})}\right)$
and hence $\lim_{n\to\infty}Y_{n}=0$ in norm. Finally, we conclude that
$\lim_{n\to\infty}S_{n}$ $=\lim_{n\to\infty}X_{n}=\lim_{n\to\infty}\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{k=0}^{m}q_{k,0}^{(m)}[m]_{q}!\lambda^{\frac{m}{2}}c(\bar{e})^{k}c^{*}(e)^{m-k}c(e)^{m}$
$=\lim_{n\to\infty}\frac{1}{n}\sum_{m=1}^{n}\sum_{k=0}^{m}q_{k,0}^{(m)}d_{m}(1-q)^{k}\lambda^{\frac{k}{2}}c(\bar{e})^{k}T_{m-k}c(e)^{k}$
where $T_{p}=(1-q)^{p}\lambda^{\frac{p}{2}}c^{*}(e)^{p}c(e)^{p}$ for any $p\geq 1$. Let $T=\lim_{p\to\infty}T_{p}$ in norm (which exists by [BMRW, Lemma 3.4] as mentioned above). Also for $0\leq k\leq m$, it is straightforward to verify that $q_{k,0}^{(m)}=\begin{pmatrix}m\\ k\end{pmatrix}_{q}$ (see for example, [BKS, Corollary 2.8]); so that for each $k\geq 0$, we have
$\lim_{m\to\infty}q_{k,0}^{(m)}=\lim_{m\to\infty}\begin{pmatrix}m\\ k\end{pmatrix}_{q}=\lim_{m\to\infty}{\frac{d_{m}}{d_{k}d_{m-k}}}=\frac{1}{d_{k}}.$
Finally, since $\lim_{m\to\infty}d_{m}=d_{\infty}$ and $\lim_{m\to\infty}T_{m-k}=T$ for all fixed $k\geq 1$, we use dominated convergence theorem to conclude that
$\lim_{m\to\infty}\sum_{k=0}^{m}q_{k,0}^{(m)}d_{m}(1-q)^{k}\lambda^{\frac{k}{2}}c(\bar{e})^{k}T_{m-k}c(e)^{k}=\sum_{k=0}^{\infty}\frac{d_{\infty}}{d_{k}}(1-q)^{k}\lambda^{\frac{k}{2}}c(\bar{e})^{k}Tc(e)^{k}.$
In particular, the Cesáro sum of the above sequence has the same limit; hence we get
$\lim_{n\to\infty}S_{n}=\lim_{n\to\infty}\frac{1}{n}\sum_{m=1}^{n}\sum_{k=0}^{m}q_{k,0}^{(m)}d_{m}(1-q)^{k}\lambda^{\frac{k}{2}}c(\bar{e})^{k}T_{m-k}c(e)^{k}=\sum_{k=0}^{\infty}\frac{d_{\infty}}{d_{k}}(1-q)^{k}\lambda^{\frac{k}{2}}c(\bar{e})^{k}Tc(e)^{k}=S.$
Finally, to prove the invertibility of $S$, we note that
$d_{\infty}^{-1}T^{-1}S-I=\sum_{k=1}^{\infty}\frac{1}{d_{k}}(1-q)^{k}\lambda^{\frac{k}{2}}T^{-1}c(\bar{e})^{k}Tc(e)^{k}$
so that
$\|d_{\infty}^{-1}T^{-1}S-I\|\leq C_{q}^{2}\sum_{k=1}^{\infty}\frac{1}{d_{k}}\lambda^{\frac{k}{2}}\|T^{-1}\|\;\|T\|$ $\leq C_{q}^{3}\|T^{-1}\|\;\|T\|\frac{\lambda^{\frac{1}{2}}}{{1-\lambda^{\frac{1}{2}}}}\leq C_{q}^{4}\frac{\lambda^{\frac{1}{2}}}{{1-\lambda^{\frac{1}{2}}}}$
where we have used the fact that $\|T^{-1}\|\;\|T\|\leq\frac{1}{d_{\infty}}\leq C_{q}$ from [BMRW, Lemma 3.4]. In particular, $d_{\infty}^{-1}T^{-1}S$ is invertible if $C_{q}^{4}\frac{\lambda^{\frac{1}{2}}}{{1-\lambda^{\frac{1}{2}}}}<1$, which is satisfied by the given value of $\lambda$ in the hypothesis. ∎
Remark 3.9.
1. Since $\|d_{\infty}^{-1}T^{-1}S-1\|\leq C_{q}^{4}\frac{\sqrt{\lambda}}{1-\sqrt{\lambda}}$, we have that
$\|d_{\infty}TS^{-1}\|=\|(I-(I-d_{\infty}^{-1}T^{-1}S))^{-1}\|\leq\frac{1}{1-\|I-d_{\infty}^{-1}T^{-1}S\|}\leq\frac{1}{1-C_{q}^{4}\frac{\sqrt{\lambda}}{1-\sqrt{\lambda}}}=\frac{1-\sqrt{\lambda}}{1-(1+C_{q}^{4})\sqrt{\lambda}}$
and since $T^{-1}\leq 2d_{\infty}^{-1}\leq 2C_{q}$ ([BMRW, Lemma 3.4]), we must have
$\|S^{-1}\|\leq\|d_{\infty}^{-1}T^{-1}\|\;\|d_{\infty}TS^{-1}\|\leq\frac{2C_{q}^{2}(1-\sqrt{\lambda})}{1-(1+C_{q}^{4})\sqrt{\lambda}}.$
2. Although we have shown that the Cesáro sum of the sequence $\{W(\bar{e}^{m}e^{m})c(\bar{e})^{m}c(e)^{m}\}_{m\geq 1}$ converges, the sequence itself does not converge. For example, this can be checked for $q=0$ where the main obstruction is that the sequence $\{\sum_{\ell=0}^{m}c(\bar{e})^{m}c(e)^{\ell}c^{*}(\bar{e})^{m-\ell}c(\bar{e})^{m}c(e)^{m}\}_{m\geq 1}$ is not convergent. This is one very important reason for us to consider Cesáro sum of such bounded sequence.
We now prove some convergence results involving both the left and right creation/annihilation operators. For this purpose, we introduce the operator $Q$ defined by
$Q=\sum_{n=0}^{\infty}q^{n}P_{n}$
where $P_{n}$ denotes the projection in $B(\mathcal{F}_{q}(\mathsf{H}_{U}))$ onto the $n$th particle space $\mathsf{H}_{U}^{\otimes_{q}^{n}}$, and the sum above converges in norm. Note that for any $f\in\mathsf{H}_{U}^{\otimes_{q}^{n}}$, we have $Qf=q^{n}f$. The following is an easy verification
(7) $r^{*}(\xi)c(\eta)=\langle\xi,\eta\rangle_{U}Q+c(\eta)r^{*}(\xi)$
for any $\xi,\eta\in\mathsf{H}_{U}$. In particular, $r^{*}(\xi)$ and $c(\eta)$ commute if $\xi$ and $\eta$ are orthogonal. Moreover, we have
(8) $Qc(\xi)=qc(\xi)Q,\quad Qr(\xi)=qr(\xi)Q$
Lemma 3.10.
For any $\xi,\eta\in\mathsf{H}_{U}$ and $m,n\in\mathbb{N}$, we have the following:
$r^{*}(\xi)^{m}c(\eta)^{n}=\sum_{k=0}^{\min\{m,n\}}\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m]_{q}!}{[m-k]_{q}!}\langle\xi,\eta\rangle_{U}^{k}c(\eta)^{n-k}Q^{k}r^{*}(\xi)^{m-k}.$
Proof.
We may and we will assume that $\langle\xi,\eta\rangle_{U}\neq 0$, otherwise there is nothing to prove since $r^{*}(\xi)c(\eta)=c(\eta)r^{*}(\xi)$ whenever $\langle\xi,\eta\rangle_{U}=0$. By replacing $\xi$ with $\frac{\xi}{\langle\eta,\xi\rangle_{U}}$ (or $\eta$ with $\frac{\eta}{\langle\xi,\eta\rangle_{U}})$, we will further assume that $\langle\xi,\eta\rangle_{U}=1$.
First, we prove the expression for $r^{*}(\xi)c(\eta)^{n}$ by induction on $n$. Indeed for $n=1$, the expression follows from (7). Assuming this to be true for $n-1$, we again use (7) and (8) to obtain
(9) $r^{*}(\xi)c(\eta)^{n}$ $=\left(r^{*}(\xi)c(\eta)\right)c(\eta)^{n-1}=Qc(\eta)^{n-1}+c(\eta)(r^{*}(\xi)c(\eta)^{n-1})$
$=q^{n-1}c(\eta)^{n-1}Q+c(\eta)\left(c(\eta)^{n-1}r^{*}(\xi)+[n-1]_{q}c(\eta)^{n-2}Q\right)$
$=c(\eta)^{n}r^{*}(\xi)+[n]_{q}c(\eta)^{n-1}Q$
as required. Now we prove the result for general $m$ and $n$ such that $n\geq m$. The case $n<m$ follows similarly. So now we fix $n$. The proof proceeds by induction on $m$. For $m=1$, the expression is proved above. By assuming the expression to be true for $m-1$, we calculate
$r^{*}(\xi)^{m}c(\eta)^{n}$ $=r^{*}(\xi)(r^{*}(\xi)^{m-1}c(\eta)^{n})$
$=r^{*}(\xi)\sum_{k=0}^{m-1}\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}c(\eta)^{n-k}Q^{k}r^{*}(\xi)^{m-1-k}$
$=\sum_{k=0}^{m-1}\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}\left(c(\eta)^{n-k}r^{*}(\xi)+[n-k]_{q}c(\eta)^{n-k-1}Q\right)Q^{k}r^{*}(\xi)^{m-1-k}$
[…]
$=\sum_{k=0}^{m-1}\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}q^{k}c(\eta)^{n-k}Q^{k}r^{*}(\xi)^{m-k}$
$\quad\quad+\sum_{k=0}^{m-1}\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}[n-k]_{q}c(\eta)^{n-k-1}Q^{k+1}r^{*}(\xi)^{m-1-k}$
$=\sum_{k=0}^{m-1}\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}q^{k}c(\eta)^{n-k}Q^{k}r^{*}(\xi)^{m-k}$
$\quad\quad+\sum_{k=1}^{m}\begin{pmatrix}n\\ k-1\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-k]_{q}!}[n-k+1]_{q}c(\eta)^{n-k}Q^{k}r^{*}(\xi)^{m-k}$
[…]
$=c(\eta)^{n}r^{*}(\xi)^{m}+\sum_{k=1}^{m-1}\left(\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}q^{k}+\begin{pmatrix}n\\ k-1\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-k]_{q}!}[n-k+1]_{q}\right)c(\eta)^{n-k}Q^{k}r^{*}(\xi)^{m-k}$
$\quad\quad\quad+\begin{pmatrix}n\\ m-1\end{pmatrix}_{q}[m-1]_{q}![n-m+1]_{q}c(\eta)^{n-m}Q^{m}$
$=\sum_{k=0}^{m}\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m]_{q}!}{[m-k]_{q}!}c(\eta)^{n-k}Q^{k}r^{*}(\xi)^{m-k}$
where to get the last expression, we have used the following calculation:
$\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}q^{k}$ $+\begin{pmatrix}n\\ k-1\end{pmatrix}_{q}\frac{[m-1]_{q}!}{[m-k]_{q}!}[n-k+1]_{q}$
$=\frac{[n]_{q}!}{[n-k]_{q}![k-1]_{q}!}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}\left(\frac{q^{k}}{[k]_{q}}+\frac{1}{[m-k]_{q}}\right)$
$=\frac{[n]_{q}!}{[n-k]_{q}![k-1]_{q}!}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}\left(\frac{q^{k}[m-k]_{q}+[k]_{q}}{[k]_{q}[m-k]_{q}}\right)$
$=\frac{[n]_{q}!}{[n-k]_{q}![k-1]_{q}!}\frac{[m-1]_{q}!}{[m-1-k]_{q}!}\left(\frac{[m]_{q}}{[k]_{q}[m-k]_{q}}\right)$
$=\begin{pmatrix}n\\ k\end{pmatrix}_{q}\frac{[m]_{q}!}{[m-k]_{q}!}.$
This completes the proof. ∎
The following proposition can be considered a stronger version of [BMRW, Lemma 3.11] in the sense of norm convergence replacing the weak convergence.
Proposition 3.11.
For any non-zero vectors $\xi,\eta\in\mathsf{H}_{U}$ with $\langle\xi,\eta\rangle_{U}=0$, we have $\lim_{n\to\infty}R_{n}=d_{\infty}^{2}P_{\Omega}$ in norm, where $R_{n}$ is the operator
$\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\|\xi\|_{U}^{-2m}\|\eta\|_{U}^{-2m}r^{*}(\xi)^{m}r^{*}(\eta)^{m}c(\xi)^{m}c(\eta)^{m}$
and $P_{\Omega}$ denotes the projection onto the vacuum space $\mathbb{C}\Omega$.
Proof.
Assume without loss of generality that $\xi,\eta$ are unit vectors. First of all, note that $r^{*}(\eta)$ and $c(\xi)$ commute as $\langle\xi,\eta\rangle_{U}=0$, so we can also write
$R_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}r^{*}(\xi)^{m}c(\xi)^{m}r^{*}(\eta)^{m}c(\eta)^{m}.$
Use Lemma 3.10 to expand $R_{n}$ as follows:
$R_{n}$ $=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\left(\sum_{k=0}^{m}\begin{pmatrix}m\\ k\end{pmatrix}_{q}\frac{[m]_{q}!}{[m-k]_{q}!}c(\xi)^{m-k}Q^{k}r^{*}(\xi)^{m-k}\right)\left(\sum_{\ell=0}^{m}\begin{pmatrix}m\\ \ell\end{pmatrix}_{q}\frac{[m]_{q}!}{[m-\ell]_{q}!}c(\eta)^{m-\ell}Q^{\ell}r^{*}(\eta)^{m-\ell}\right)$
$=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{k,\ell=0}^{m}\begin{pmatrix}m\\ k\end{pmatrix}_{q}\begin{pmatrix}m\\ \ell\end{pmatrix}_{q}\frac{[m]_{q}!^{2}}{[m-k]_{q}![m-\ell]_{q}!}c(\xi)^{m-k}Q^{k}c(\eta)^{m-\ell}r^{*}(\xi)^{m-k}Q^{\ell}r^{*}(\eta)^{m-\ell}$
(since $\hskip 289.07999pt(\text{since }r^{*}(\xi)c(\eta)=c(\eta)r^{*}(\xi))$
$=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{k,\ell=0}^{m}q^{k(m-\ell)+\ell(m-k)}\begin{pmatrix}m\\ k\end{pmatrix}_{q}\begin{pmatrix}m\\ \ell\end{pmatrix}_{q}\frac{[m]_{q}!^{2}}{[m-k]_{q}![m-\ell]_{q}!}c(\xi)^{m-k}c(\eta)^{m-\ell}Q^{k+\ell}r^{*}(\xi)^{m-k}r^{*}(\eta)^{m-\ell}$
[…]
$=\frac{1}{n}\sum_{m=1}^{n}d_{m}^{2}Q^{2m}+\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}c(\xi)^{m}c(\eta)^{m}r^{*}(\xi)^{m}r^{*}(\eta)^{m}+X_{n}$
where the sum is divided over $(k,\ell=m)$, $(k,\ell=0)$ and the rest, i.e. $X_{n}$ is the operator
$X_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\sum_{\underset{0<k+\ell<2m}{k,\ell=0}}^{m}q^{m(k+\ell)-2k\ell}\begin{pmatrix}m\\ k\end{pmatrix}_{q}\begin{pmatrix}m\\ \ell\end{pmatrix}_{q}\frac{[m]_{q}!^{2}}{[m-k]_{q}![m-\ell]_{q}!}c(\xi)^{m-k}c(\eta)^{m-\ell}Q^{k+\ell}r^{*}(\xi)^{m-k}r^{*}(\eta)^{m-\ell}.$
Thus we have divided $R_{n}$ as a sum of three terms, and we will calculate norm limit of each term. To get the limit of the first term, first verify that $Q^{2m}$ converges to $P_{\Omega}$ as $m\to\infty$, and therefore the average $\frac{1}{n}\sum_{m=1}^{n}d_{m}^{2}Q^{2m}$ converges to $d_{\infty}^{2}P_{\Omega}$ in norm. The second term
$\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}c(\xi)^{m}c(\eta)^{m}r^{*}(\xi)^{m}r^{*}(\eta)^{m}$
[…]
$\|X_{n}\|$ $\leq\frac{C_{q}^{2}}{n}\sum_{m=1}^{n}|q|^{m}(1-q)^{2m}\sum_{\underset{0<k+\ell<2m}{k,\ell=0}}^{m}\frac{[m]_{q}!^{2}}{[k]_{q}![\ell]_{q}![m-k]_{q}![m-\ell]_{q}!}\frac{[m]_{q}!^{2}}{[m-k]_{q}![m-\ell]_{q}!}\left([m-k]_{q}![m-\ell]_{q}!\right)$
$\leq\frac{C_{q}^{2}}{n}\sum_{m=1}^{n}|q|^{m}\sum_{{k,\ell=0}}^{m}\frac{d_{m}^{4}}{d_{k}d_{\ell}d_{m-k}d_{m-\ell}}$
$\leq\frac{C_{q}^{10}}{n}\sum_{m=1}^{n}\sum_{{k,\ell=0}}^{m}|q|^{m}$
$\leq\frac{C_{q}^{10}}{n}\sum_{m=1}^{n}m^{2}|q|^{m}.$
The sequence $\{m^{2}q^{m}\}_{m\geq 1}$ converges to 00 as $m\to\infty$, and so does its average; hence we conclude that $X_{n}$ converges to 00 in norm as $n\to\infty$. ∎
We end this section by giving an expression for the right Wick product $W_{r}(\bar{e}^{n}e^{n})$, which can easily be proved by utilizing Proposition 2.2.
Corollary 3.12.
For $e,\bar{e}$ and $\lambda$ as before, the right Wick product $W_{r}(\bar{e}^{n}e^{n})$ is given as
$W_{r}(\bar{e}^{n}e^{n})$ $=\sum_{k,\ell=0}^{n}q^{(n)}_{k,\ell}r(e)^{k}r(\bar{e})^{\ell}r^{*}(\bar{e}^{r})^{n-k}r^{*}(\overline{\bar{e}}^{r})^{n-\ell}$
$=\sum_{k,\ell=0}^{n}q^{(n)}_{k,\ell}\lambda^{k-\ell}r(e)^{k}r(\bar{e})^{\ell}r^{*}(\bar{e})^{n-k}r^{*}(e)^{n-\ell}$
where $q_{k,\ell}^{(m)}$ are the same scalars as in the proof of Lemma 3.6.
4. Main result
To establish the fullness of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$, we rely on the equivalent condition of the triviality of the relative commutant within the ultraproduct. We briefly recall the construction of the von Neumann ultraproduct. Let $\mathsf{M}$ be a von Neumann algebra equipped with a normal faithful state $\varphi$, and let $\omega$ be a free ultrafilter over the natural numbers. Denote by $\prod_{n\geq 1}\mathsf{M}$ the algebra of all bounded sequences $(x_{n})$ with each $x_{n}\in\mathsf{M}$. Let $I_{\omega}$ be the two sided ideal of $\prod_{n\geq 1}\mathsf{M}$ given by
$I_{\omega}=\left\{(x_{n})\in\prod_{n\geq 1}\mathsf{M};\lim_{n\to\infty}\varphi(x_{n}^{*}x_{n}+x_{n}x_{n}^{*})=0\right\}.$
The quotient $\mathsf{M}^{\omega}:=\prod_{n\geq 1}\mathsf{M}/I_{\omega}$ is a well-defined von Neumann algebra, called ultraproduct of $\mathsf{M}$. A typical element in $\mathsf{M}^{\omega}$ will be denoted as $(x_{n})_{\omega}$ which is the equivalence class of the sequence $(x_{n})$ in the quotient. It is known that $\mathsf{M}$ is a full factor if and only if $\mathsf{M}^{\prime}\cap\mathsf{M}^{\omega}=\mathbb{C}$ [Con1]. The proof of our main result on fullness uses this particular equivalent condition. In what follows, one can also directly employ the equivalent condition of fullness as formulated in [Mar, Corollary B].
Proof of Theorem 1.1. If $\mathsf{H}_{\mathbb{R}}$ is finite-dimensional, then $\mathsf{M}_{q}$ is full by [KSW, Theorem 2.7]. If $\mathsf{H}_{\mathbb{R}}$ has a non-zero weakly mixing part or if the point spectrum $\sigma_{p}(A)$ of eigenvalues of the analytic generator $A$ of $\{U_{t}\}_{t\in\mathbb{R}}$ is infinite and bounded, then $\mathsf{M}_{q}$ is full by [HI, Theorem 6.2]. So assume that $\mathsf{H}_{\mathbb{R}}$ is infinite-dimensional and there is a strictly decreasing sequence of eigenvalues of $A$ converging to 00.
Now let $\omega$ be a fixed non-principal ultrafilter. Let $(X_{k})_{\omega}\in\mathsf{M}_{q}^{\prime}\cap\mathsf{M}_{q}^{\omega}$. By replacing each $X_{k}$ by $X_{k}-\varphi(X_{k})$, we may assume that $\varphi(X_{k})=0$ for all $k\geq 1$. In particular, we have $X_{k}\Omega,X_{k}^{*}\Omega\in\mathcal{F}_{q}(\mathsf{H}_{U})\ominus\mathbb{C}\Omega$ for all $k\geq 1$. We will show that $(X_{k})_{\omega}=0$ in $\mathsf{M}_{q}^{\omega}$ i.e. $\lim_{k\to\omega}\|X_{k}\Omega\|_{q}=0=\lim_{k\to\omega}\|X_{k}^{*}\Omega\|_{q}$.
Fix $\lambda\in\sigma_{p}(A)$ with $10<\lambda<$ 1, and let $e\in\mathsf{H}_{\mathbb{C}}$ be such that $Ae=\lambda^{-1}e$ and $A\bar{e}=\lambda\bar{e}$ with $\|e\|_{U}=\lambda^{-\frac{1}{4}}$ and $\|\bar{e}\|_{U}=\lambda^{\frac{1}{4}}$. Consider the operator $S_{n}$
$S_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}W(\bar{e}^{m}e^{m})c(\bar{e})^{m}c(e)^{m}.$
Assume that $\lambda<(1+C_{q}^{4})^{-2}$ so that the norm limit $S$ of $S_{n}$ exists and is invertible by Proposition 3.8. We observe
$\lim_{k\to\omega}\|X_{k}\Omega\|_{q}^{2}$ $=\lim_{k\to\omega}\langle X_{k}\Omega,X_{k}\Omega\rangle_{q}=\lim_{k\to\omega}\lim_{n\to\infty}\langle X_{k}\Omega,S_{n}S^{-1}X_{k}\Omega\rangle_{q}=\lim_{n\to\infty}\lim_{k\to\omega}\langle X_{k}\Omega,S_{n}S^{-1}X_{k}\Omega\rangle_{q}$
where we could exchange the limit because $S_{n}S^{-1}$ converges in norm to identity as $n\to\infty$. Now we use the expressions $XX_{k}-X_{k}X\to 0$-strongly as $k\to\omega$ for all $X\in\mathsf{M}_{q}$, and $X_{k}X^{\prime}-X^{\prime}X_{k}=0$ for all $X^{\prime}\in\mathsf{M}_{q}^{\prime}$ to get
$\lim_{k\to\omega}\|X_{k}\Omega\|_{q}^{2}$ $=\lim_{n\to\infty}\lim_{k\to\omega}\langle X_{k}\Omega,S_{n}S^{-1}X_{k}\Omega\rangle_{q}$
$=\lim_{n\to\infty}\lim_{k\to\omega}\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\langle X_{k}\Omega,W(\bar{e}^{m}e^{m})c(\bar{e})^{m}c(e)^{m}S^{-1}X_{k}\Omega\rangle_{q}$
$=\lim_{n\to\infty}\lim_{k\to\omega}\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\langle W(\bar{e}^{m}e^{m})X_{k}\Omega,c(\bar{e})^{m}c(e)^{m}S^{-1}X_{k}\Omega\rangle_{q}$
$=\lim_{n\to\infty}\lim_{k\to\omega}\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\langle X_{k}W(\bar{e}^{m}e^{m})\Omega,c(\bar{e})^{m}c(e)^{m}S^{-1}X_{k}\Omega\rangle_{q}$
$=\lim_{n\to\infty}\lim_{k\to\omega}\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\langle X_{k}W_{r}(\bar{e}^{m}e^{m})\Omega,c(\bar{e})^{m}c(e)^{m}S^{-1}X_{n}\Omega\rangle_{q}$
$=\lim_{n\to\infty}\lim_{k\to\omega}\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\langle X_{k}\Omega,W_{r}(\bar{e}^{m}e^{m})c(\bar{e})^{m}c(e)^{m}S^{-1}X_{k}\Omega\rangle_{q}$
$=\lim_{n\to\infty}\lim_{k\to\omega}\langle X_{k}\Omega,T_{n}S^{-1}X_{k}\Omega\rangle_{q}+\lim_{n\to\infty}\lim_{k\to\omega}\langle X_{k}\Omega,R_{n}S^{-1}X_{k}\Omega\rangle_{q}$
where $T_{n}$ and $R_{n}$ are the following operators:
$T_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}\left(W_{r}(\bar{e}^{m}e^{m})-r^{*}(\bar{e})^{m}r^{*}(e)^{m}-r(e)^{m}r(\bar{e})^{m}\right)c(\bar{e})^{m}c(e)^{m}$
$R_{n}=\frac{1}{n}\sum_{m=1}^{n}(1-q)^{2m}(r^{*}(\bar{e})^{m}r^{*}(e)^{m}+r(e)^{m}r(\bar{e})^{m})c(\bar{e})^{m}c(e)^{m}.$
Propositions 3.4 and 3.11 tell us that the operators $R_{n}$ (or $R_{n}^{*}$) when restricted to $\mathcal{F}_{q}(\mathsf{H}_{U})\ominus\mathbb{C}\Omega$ converge to 00 in norm as $n\to\infty$. Since $X_{k}\Omega\in\mathcal{F}_{q}(\mathsf{H}_{U})\ominus\mathbb{C}\Omega$, it follows that
$\lim_{n\to\infty}\lim_{k\to\omega}\langle X_{k}\Omega,R_{n}S^{-1}X_{k}\Omega\rangle_{q}=0.$
Our next aim is to show that the operators $T_{n}$ are bounded by $\frac{2C_{q}^{6}\sqrt{\lambda}}{1-\sqrt{\lambda}}$ as $n\to\infty$. First write using Corollary 3.12 the following:
$W_{r}(\bar{e}^{m}e^{m})=\sum_{k,\ell=0}^{m}q^{(m)}_{k,\ell}\lambda^{k-\ell}r(e)^{k}r(\bar{e})^{\ell}r^{*}(\bar{e})^{m-k}r^{*}(e)^{m-\ell}.$
and then note that
$W_{r}(\bar{e}^{m}e^{m})$ $-r^{*}(\bar{e})^{m}r^{*}(e)^{m}-r(e)^{m}r(\bar{e})^{m}=\sum_{\ell=1}^{m}\sum_{k=0}^{m-1}q^{(m)}_{k,\ell}\lambda^{k-\ell}r(e)^{k}r(\bar{e})^{\ell}r^{*}(\bar{e})^{m-k}r^{*}(e)^{m-\ell}$
$+\sum_{\ell=0}^{m-1}q_{m,\ell}^{(m)}\lambda^{m-\ell}r(e)^{m}r(\bar{e})^{\ell}r^{*}(e)^{m-\ell}+\sum_{k=1}^{m}q_{k,0}^{(m)}\lambda^{k}r(e)^{k}r^{*}(\bar{e})^{m-k}r^{*}(e)^{m}$
so that the upper bound for $T_{n}$ can be given as follows:
$\|T_{n}\|$ $\leq\frac{C_{q}^{6}}{n}\sum_{m=1}^{n}\left(\sum_{\ell=1}^{m}\sum_{k=0}^{m-1}|q|^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}+\sum_{\ell=0}^{m-1}\lambda^{\frac{m-\ell}{2}}+\sum_{k=1}^{m}\lambda^{\frac{k}{2}}\right)\leq{C_{q}^{6}}s_{n}+\frac{2C_{q}^{6}\sqrt{\lambda}}{1-\sqrt{\lambda}}.$
where $s_{n}$ is the sequence as in Lemma 3.7, which converges to 00 by the same Lemma. Thus, we have shown that
$\lim_{k\to\omega}\|X_{k}\Omega\|_{q}^{2}$ $=\lim_{n\to\infty}\lim_{k\to\omega}\langle X_{k}\Omega,T_{n}S^{-1}X_{k}\Omega\rangle_{q}$
$\leq\left(\limsup_{n\to\infty}\|T_{n}\|\right)\;\|S^{-1}\|\;L$
$\leq\frac{2LC_{q}^{6}\sqrt{\lambda}}{1-\sqrt{\lambda}}\cdot\frac{2C_{q}^{2}(1-\sqrt{\lambda})}{1-(1+C_{q}^{4})\sqrt{\lambda}}$
where $L=\sup_{k\geq 1}\|X_{k}\Omega\|^{2}$, and we have used the upper bound for $\|S^{-1}\|$ from Remark 3.9. By our assumption, we have a sequence of such $\lambda$’s which converges to 00; hence it follows that $\lim_{k\to\omega}\|X_{k}\Omega\|_{q}=0$. A similar argument shows that $\lim_{k\to\omega}\|X_{k}^{*}\Omega\|_{q}=0$. This completes the proof.
Acknowledgments. We are grateful to Adam Skalski and Zhiyuan Yang for going through the initial draft, and providing valuable suggestions; particularly, Zhiyuan suggested a shortened proof of Proposition 3.11 by utilizing the operator $Q$. We thank the referee for their careful reading of the manuscript and for several useful comments. This work was partially completed when M.K. was visiting S.W. at Harbin Institute of Technology in Aug-Sep 2023 for a month. M.K. extends appreciation to HIT for their warm hospitality and to S.W. for the invitation and financial support.
M.K. was partially supported by the National Science Center (NCN) grant no. 2020/39/I/ST1/01566. The project is co-financed by the Polish National Agency for Academic Exchange within Polish Returns Program for Mateusz Wasilewski. S.W. was partially supported by the NSF of China (No.12031004, No.12301161) and the Fundamental Research Funds for the Central Universities.